If a box of sweets is divided among 24 children, they will get 5 sweets each. How many would each get, if the number of the children is reduced by 4?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Each child gets = 5 sweets
24 children will get 24×5 = 120 sweets.

Total number of sweets = 120
If the number of children is reduced by 4, then children left = 24-4 = 20

Now each child will get sweets = 120/20 = 6 sweets

Thanks